can you show me step by step how to get solve this? thank you
Question 1 Solve the following circuit using superposition Is S R₂ R₃ < 1.2 k2 18 12 600 MA
Find Vo using superposition. SR 512 200 mA 330
Answer #1

Summary- This problem is solved by using the basics of Network Analysis using the superposition theorem. Below are the details of solution uploaded in sequence.

10 & w NW To ww polalled = 5x3 573 Step G Reduce Sources the circuit as possible without disturbing the and then Apply superNow, circuit will be. 3 V 4 ww Fig ③ :- ckt De to source (Vs, =lo v alone) node Apply Nodal Analysis at V By KCL Va + Va t xNow, using Voltage Division Rue, E-376 * 42 (RAR) = -1-376 * 4 = 85 . (4+1) (s. Due to opposite terminal) Voltage Division RuStep ③ - NOW Due to let Now, circuit will be current source alone v be voltage Pource Tranformation Technique. VU + Fig. © CStep 4 - Now, Voltage source (V3 =10) alone... Circuit will be 3x4 (3 . W 7 Nu + w alin Fig. ckt Due to Vss alone Reduced OktBy super position theorem, Vo = V + Vor at von (Due to (Due to (Due to Vsi alone) Is alone) Viss alone) from, egnis 6 § 256 8